OpenStudy (anonymous):

Triangle ABC is congruent to Triangle PQR. can you tell me the value of 'y'?

OpenStudy (anonymous):



OpenStudy (anonymous):

CPCTC

OpenStudy (anonymous):

Corrresponding Parts of Congruent Triangles are Congruent Can you identify what corresponding sides are congruent?

OpenStudy (anonymous):

corresponding sides of congruent triangle are proportional

OpenStudy (anonymous):

@stupidinmath still there?

OpenStudy (anonymous):

@giligails are you referring to similar triangles?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the two triangles are conguent, given

OpenStudy (anonymous):

@stupidinmath still there? yuhoo

OpenStudy (akashdeepdeb):

Is this query cleared? :)

OpenStudy (anonymous):

CB is congruent to QR x+9 is congruent to 20 2y-1 is congruent to y+8

OpenStudy (anonymous):

@stupidinmath yes that's it, then you can solve the value of y now, just equate it.

OpenStudy (anonymous):

its hard when idk the other side of the triangle :| like CB & QR

OpenStudy (anonymous):

why bother CB and QR, you are just ask to solve the value of y, and you just said earlier that "2y-1 is congruent to y+8", from here you can solve y.

OpenStudy (anonymous):

is y=9 correct?

OpenStudy (anonymous):

@Jaseem18

OpenStudy (anonymous):

Yes, you got it.

OpenStudy (anonymous):

thank you @Jaseem18 :)

OpenStudy (phi):

2y-1 is congruent to y+8 means the length of both sides are the same in other words, you can say 2y-1 = y+8 add -y to both sides 2y - y -1 = y+ -y + 8 y + -y is 0 on the right side. 2 y's take away 1 y leaves 1 y on the left side 1y - 1 = 8 now add +1 to both sides y -1 +1= 8+1 y=9

OpenStudy (anonymous):

thank you @phi :D

OpenStudy (anonymous):

you are welcome